Fermi National Laboratory


Accelerator Update

Wednesday November 12
At 8:16 AM Operations had a problem with the readbacks for the water pumps at Tevatron (TeV) sector A1. Operators power cycled the MADC.

Water technicians began a hydraulic test on two LCW systems in the Main Injector (MI) at 9:00 AM. The test lasted all day.

A Linac expert switched over to the H- Source at 1 PM.

Operations received written permission to deliver beam to the Antiproton Source (Pbar) at 4 PM.

Operators began interlocking the Collider Detector Facility (CDF) and D-zero (D0) at 4 PM, and they interlocked MI at 9 PM.

At 6 PM ComEd called and informed Fermilab that there was a problem off site that required them to cut power to the Kautz Rd substation. Experts switched high voltage for important systems to other sources.

Thursday November 13
At 6 AM operators discovered a problem with the TeV sector C2 dump. A breaker reset resolved the problem.

Early in the morning, high voltage experts discovered a static line that had come loose during the night's heavy winds. After contacting ComEd a site wide power outage was planned for noon and then switched to early evening.

Experts turned off Linac and Preacc for a proposed power outage at noon, which never occurred. Experts decide to keep systems off until power returned to its normal configuration on Friday.

The Fermi site wide power outage occurred at 5 PM to switch power from the Master substation to the Kautz Rd substation. Power was restored thirteen minutes later. Experts and techs began restoring systems. High voltage experts called in an outside contractor to repair the static line. They had the static line repaired by 10:30 PM.

Operators successfully hipotted the TeV and MI before the end of the evening shift.

Friday November 14
The midnight shift continued to recover systems.

Future Plans
The shutdown is over. We had hoped to be running beam for tune up and studies by this morning, but the Kautz Rd power outage for ComEd and then our own Master substation outage for the static line repair have delayed this by two to three days primarily due to TeV cryo systems that need to be cooled back down. Although this is unfortunate, if either of these problems had caused an outage that we hadn't prepared for, the delay would have been much longer.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 11/14/2003   email Fermilab